OpenStudy (anonymous):

It takes 22 minutes for a certain bacteria population to triple. At 8:00 am, the bacteria count was 3,000,000. What is the best estimate (in millions of units) of the population at 9:06 am on the same morning.

OpenStudy (tkhunny):

How many time periods of 22 minutes have passed?

OpenStudy (anonymous):

I don't know

OpenStudy (tkhunny):

NonResponsive. Start at 8:00 AM and end at 9:06 AM. You CAN answer my question.

OpenStudy (anonymous):

3 times (i think)

OpenStudy (tkhunny):

Think or know. You have to be sure. Prove it. Show how you did it.

OpenStudy (anonymous):

I know because 22*3=66 minutes and that= 1hour and 6 minutes

OpenStudy (tkhunny):

That's better. Let's try a simple example. If you start with 2 things, what happens when we TRIPLE it three times?

OpenStudy (anonymous):

it=18

OpenStudy (anonymous):

yes i agree

OpenStudy (tkhunny):

Sorry, I fixed a typo. 2*3 = 6 6*3 = 18 18*3 = 54 You tripled it only twice. Agreed?

OpenStudy (tkhunny):

Okay, now to your problem. You have 3 things. Triple it three times. What do you get?

OpenStudy (anonymous):

ohh so it would be 27,000,000

OpenStudy (tkhunny):

No. That was only twice. 3*3 = 9 9*3 = 27 27*3 = 81 Something in your head gets so excited at 2 that you forget to do #3!! Good work (except for counting). :-)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

!!